# Homological stratification and descent

### v1

@@ 11. Equivariant applications
For foundational material on $\infty$-categories of equivariant $G$-spectra, we refer the reader to [MNN17, BGH20, BCH+24a], which are based on the more classical references [LMS86, May96, MM02]. We write $\operatorname{Sp}_{G}$ for this category. Recall that this is a presentably symmetric monoidal stable $\infty$-category which is rigidly-compactly generated by the orbits $G/H_{+}$ associated to the (conjugacy classes) of closed subgroups $H\leq G$. Here, we omit the suspension spectrum from our notation.
@@
induced by base-change $S_{G}^{0}\to R$ and the $H$-geometric fixed point functors for $\operatorname{Mod}_{G}(R)$ and $\operatorname{Sp}_{G}$. This diagram shows that the images of the maps $\varphi^{h}_{H}$ are disjoint in $\operatorname{Spc}^{h}(\operatorname{Mod}_{G}(R)^{c})$ for non-conjugate subgroups, since this is true for the $R=S^{0}_{G}$ case just mentioned. As a consequence of 11.5 and 7.8, these images also cover the homological spectrum, which gives the desired decomposition of the homological spectrum. The analogous statement for the tt-spectrum is proved similarly. ∎
@@
{localizing ideals of ModG(R)}{subsets of $\big{\{}\text{localizing ideals of $\operatorname{Mod}_{G}(R)$}\big{\}}\xrightarrow{\cong}\big{\{}\text{subsets of $\bigsqcup_{H\in\operatorname{Sub}(G)/G}\operatorname{im}\varphi^{h}_{H}$}\big{\}}.$
@@
The first part follows from 7.21, which applies due to 11.5. The claimed bijection is then 3.1 combined with 11.6. ∎
@@
We next make the previous proposition more explicit in the special case of commutative equivariant ring spectra with trivial action, i.e., those that arise via inflation from (non-equivariant) commutative ring spectra. This provides a generalization of [BHS23b, Theorem 15.1], extending it from finite groups to compact Lie groups and removing any finiteness assumptions on the spectrum.
@@ 12. Open questions
For a field $k$, the spectrum of the ring $k[x_{1},x_{2},\ldots]$ is not weakly noetherian; this follows, for example, from [BCHS23, Proposition 4.13]. Even for this relatively simple non-noetherian ring, we are unable to determine whether (*12.3) holds.
@@
As noted above, for an object $t\in\mathscr{T}$, the Balmer–Favi–Sanders support $\operatorname{Supp}(t)$ is always closed in the localizing topology on $\operatorname{Spc}(\mathscr{T}^{c})$. On the other hand, since $\operatorname{Supp}^{h}(\coprod_{\mathscr{B}\in S}E_{\mathscr{B}})=S$ for any subset $S\subseteq\operatorname{Spc}(\mathscr{T}^{c})$, the homological support $\operatorname{Supp}^{h}(t)$ takes values in arbitrary subsets of $\operatorname{Spc}^{h}(\mathscr{T}^{c})$. This complicates the relationship between $\pi(\operatorname{Supp}^{h}(t))$ and $\operatorname{Supp}(t)$. Nonetheless, one has the following strengthening of 5.6; for unfamiliar terminology, we refer the reader to [Zou23].
@@
We always have $\pi(\operatorname{Supp}^{h}(t))\subseteq\operatorname{Supp}(t)$ by [Zou23, Lemma 8.5] and the latter is localizing closed, hence $\overline{\pi(\operatorname{Supp}^{h}(t))}^{\mathrm{loc}}\subseteq\operatorname{Supp}(t)$. For the reverse inclusion, recall that weakly visible subsets form a basis of open subsets for the localizing topology. If $\mathscr{P}\notin\overline{\pi(\operatorname{Supp}^{h}(t))}^{\mathrm{loc}}$ then there exists a weakly visible subset $W\ni\mathscr{P}$ such that $\pi(\operatorname{Supp}^{h}(t))\cap W=\varnothing$ and hence $\pi^{-1}(W)\cap\operatorname{Supp}^{h}(t)=\varnothing$. By the tensor product formula and 5.3, we obtain $\operatorname{Supp}^{h}(t\otimes g_{W})=\varnothing$. It then follows from h-detection that $t\otimes g_{W}=0$. Therefore, $\mathscr{P}\not\in\operatorname{Supp}^{h}(t)$. ∎
@@
One can therefore ask the following a priori weaker variant of 12.9:
Question 12.12.
@@
We have introduced a notion of homological cosupport and proved in 5.5 that

### v2

@@ 11. Equivariant applications
For foundational material on $\infty$-categories of equivariant $G$-spectra, we refer the reader to [MNN17, BGH20, BCH+25], which are based on the more classical references [LMS86, MAY96, MM02]. We write $\operatorname{Sp}_{G}$ for this category. Recall that this is a presentably symmetric monoidal stable $\infty$-category which is rigidly-compactly generated by the orbits $G/H_{+}$ associated to the (conjugacy classes) of closed subgroups $H\leq G$. Here, we omit the suspension spectrum from our notation.
@@
induced by base-change $S_{G}^{0}\to R$ and the $H$-geometric fixed point functors for $\operatorname{Mod}_{G}(R)$ and $\operatorname{Sp}_{G}$. This diagram shows that the images of the maps $\varphi^{h}_{H}$ are disjoint in $\operatorname{Spc}^{h}(\operatorname{Mod}_{G}(R)^{c})$ for non-conjugate subgroups, since this is true for the $R=S^{0}_{G}$ case just mentioned. As a consequence of 11.5 and 7.9, these images also cover the homological spectrum, which gives the desired decomposition of the homological spectrum. The analogous statement for the tt-spectrum is proved similarly. ∎
@@
$\big\{\text{localizing ideals of $\operatorname{Mod}_{G}(R)$}\big\}\xrightarrow{\cong}\big\{\text{subsets of $\bigsqcup_{H\in\operatorname{Sub}(G)/G}\operatorname{im}\varphi^{h}_{H}$}\big\}.$
@@
The first part follows from 7.22, which applies due to 11.5. The claimed bijection is then 3.1 combined with 11.6. ∎
@@
We next make the previous proposition more explicit in the special case of commutative equivariant ring spectra with trivial action, i.e., those that arise via inflation from (non-equivariant) commutative ring spectra. This provides a generalization of [BHS23d, Theorem 15.1], extending it from finite groups to compact Lie groups and removing any topological assumptions on the spectrum.
@@ 12. Open questions
For a field $k$, the spectrum of the ring $k[x_{1},x_{2},\ldots]$ is not weakly noetherian; this follows, for example, from [BCH+23b, Proposition 4.13]. Even for this relatively simple non-noetherian ring, we are unable to determine whether (*12.3) holds.
@@
As noted above, for an object $t\in\mathscr{T}$, the Balmer–Favi–Sanders support $\operatorname{Supp}(t)$ is always closed in the localizing topology on $\operatorname{Spc}(\mathscr{T}^{c})$. On the other hand, since $\operatorname{Supp}^{h}(\coprod_{\mathscr{B}\in S}E_{\mathscr{B}})=S$ for any subset $S\subseteq\operatorname{Spc}(\mathscr{T}^{c})$, the homological support $\operatorname{Supp}^{h}(t)$ takes values in arbitrary subsets of $\operatorname{Spc}^{h}(\mathscr{T}^{c})$. This complicates the relationship between $\pi(\operatorname{Supp}^{h}(t))$ and $\operatorname{Supp}(t)$. Nonetheless, one has the following strengthening of 5.7; for unfamiliar terminology, we refer the reader to [ZOU23].
@@
We always have $\pi(\operatorname{Supp}^{h}(t))\subseteq\operatorname{Supp}(t)$ by [ZOU23, Lemma 8.5] and the latter is localizing closed, hence $\overline{\pi(\operatorname{Supp}^{h}(t))}^{\mathrm{loc}}\subseteq\operatorname{Supp}(t)$. For the reverse inclusion, recall that weakly visible subsets form a basis of open subsets for the localizing topology. If $\mathscr{P}\notin\overline{\pi(\operatorname{Supp}^{h}(t))}^{\mathrm{loc}}$ then there exists a weakly visible subset $W\ni\mathscr{P}$ such that $\pi(\operatorname{Supp}^{h}(t))\cap W=\varnothing$ and hence $\pi^{-1}(W)\cap\operatorname{Supp}^{h}(t)=\varnothing$. By the tensor product formula and 5.4, we obtain $\operatorname{Supp}^{h}(t\otimes g_{W})=\varnothing$. It then follows from h-detection that $t\otimes g_{W}=0$. Therefore, $\mathscr{P}\not\in\operatorname{Supp}(t)$ by the definition of the Balmer–Favi–Sanders support; see [ZOU23, Definition 5.17]. ∎
12.8.
We also have the following strengthening of 5.11:
Proposition 12.9.
For any weak ring $w\in\mathscr{T}$, we have
$\overline{\pi(\operatorname{Supp}^{h}(w))}^{\mathrm{loc}}=\operatorname{Supp}(w).$
Proof.
The $\subseteq$ inclusion always holds, as explained in the proof of 12.7. Moreover, as before, if $\mathscr{P}\not\in\overline{\pi(\operatorname{Supp}^{h}(w))}^{\mathrm{loc}}$ then there exists a weakly visible subset $W\ni\mathscr{P}$ such that $\pi^{-1}(W)\cap\operatorname{Supp}^{h}(w)=\varnothing$ and it suffices to establish that $g_{W}\otimes w=0$}. Write $W=Y_{1}\cap Y_{2}^{c}$. Since $Y_{1}$ is a union of Thomason closed sets, we may assume without loss of generality that $Y_{1}$ is a Thomason closed set. Then $Y_{1}=\operatorname{supp}(x)$ for some $x\in\mathscr{T}^{c}$ and replacing $x$ by $x\otimes x^{\vee}$ we may assume that $x$ is a (weak) ring. Observe then that
$\pi^{-1}(W)=\pi^{-1}(\operatorname{supp}(x))\cap\pi^{-1}(Y_{2}^{c})=\operatorname{Supp}^{h}(x)\cap\operatorname{Supp}^{h}(f_{Y_{2}})$
by 5.2 and 5.4. Hence
$\varnothing=\operatorname{Supp}^{h}(x)\cap\operatorname{Supp}^{h}(f_{Y_{2}})\cap\operatorname{Supp}^{h}(w)=\operatorname{Supp}^{h}(x\otimes f_{Y_{2}}\otimes w).$
This implies that $x\otimes f_{Y_{2}}\otimes w=0$ since the homological support has the detection property for weak rings (1.8). We conclude that $e_{\operatorname{supp}(x)}\otimes f_{Y_{2}}\otimes w=0$. That is, $g_{W}\otimes x=0$, as desired. ∎
@@
One can therefore ask the following a priori weaker variant of 12.11:
Question 12.14.
@@
We have introduced a notion of homological cosupport and proved in 5.6 that